# Preprint Notice

### v3

@@ Abstract
I Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps [7]. However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
@@ I Introduction
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap manifests as lower than desired data rates, even on 10 Gbps links and modest commodity hardware. Higher-speed networks only amplify its effect. The “Drainage Basin Pattern” conceptual model (Fig. 1) addresses it by shifting the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment). It also enables efficient 1–10 Gbps transfers at the edge. Detailed methodology and measurement results appear in Sections III and IV, where we analyze how host configurations influence end-to-end throughput across multiple deployment scenarios. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
@@
Figure 1: The experience of moving data for most practitioners is typically limited to the source of the river. This corresponds to end-user activities, such as transferring photos and videos from mobile phones or moving spreadsheets from one folder to another. At such low data rates, operations appear simple. This may explain why efficient data transfer has historically received limited attention, motivating the introduction of the Drainage Basin Pattern.
@@
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II [53], sustained petabyte transfers [14, 15], and international deployments [28]—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations [34, 33, 32], historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) [52]. This cross-generational validation confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
@@
A key architectural component for achieving predictable performance is the use of high-speed burst buffers [5] as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) [40] and official ESnet evaluations [34, 33, 32], underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
@@
The core contribution of this paper is the identification and validation of six widely held engineering paradigms. They continue to shape expectations, system design, and investment decisions in production end-to-end data movement. Through production-scale deployments and externally evaluated operational trials, we show why they often fail to predict achievable performance across the full edge-to-core spectrum. Reliable and predictable data movement emerges from holistic, co-designed, and tightly integrated hardware-software systems, instead of isolated optimization of individual components. The end-to-end behavior of these systems can be reasoned about using the proposed Drainage Basin Pattern.
@@
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the emphasis placed on standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for high-performance CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 [45] or free rebuilds such as Rocky Linux 9.6 [49].
II Background / Related Work
Research programs such as the Global Research Platform (GRP) [54] have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop [55] and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem [17, 21].
@@ II Background / Related Work
II-A Burst Buffers and Data Staging
We now introduce the burst buffer [5]—a high-speed intermediate storage layer. Originally pioneered in supercomputing to absorb data for high-performance computing (HPC) nodes at extreme velocities, we generalize this concept for sustained, wide-area data movement.
@@ II-A Burst Buffers and Data Staging
The burst buffer serves as a staging area between production storage and the network. A vendor-agnostic data movement appliance design employs a unified software data mover (Zettar zx [61]. See Table I for a summary of its features and capabilities) and burst buffers to create a predictable performance envelope that enables sustained near-line-rate data movement between burst buffers (fast ($100~\mathrm{Gbps}$) and high-speed ($\geq 100~\mathrm{Gbps}$) WAN links)
In this architecture, the burst buffer serves not only as a fast storage tier, but as a deliberate decoupling mechanism between the data mover and production storage services that are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors.
@@
II-B Bulk and Streaming Transfers
@@
  • Streaming Transfer: the movement of a dynamic, actively growing dataset where data is transferred concurrently as it is being generated and written to the source storage. This type is essential for real-time workloads, such as data acquisition from scientific instruments like LCLS-II [53] and large-scale bioscience imaging operations (e.g., Cryo-EM, virtual staining) that, due to their inherent data rates, will require such concurrent, line-rate data movement to achieve efficient production workflows.
This distinction is critical because streaming transfers demand both high storage throughput and low storage latency, stressing storage subsystems, memory hierarchies, and scheduling mechanisms in ways that differ fundamentally from bulk transfer. As a result, streaming workloads require architectures designed for sustained concurrency and flow rather than peak performance under isolated benchmarks.
@@
The architecture presented in this work eliminates this source of variability by removing centralized orchestration from the data path entirely. Data movement is executed as a decentralized, peer-to-peer service, coordinated implicitly through asynchronous buffer state rather than explicit global scheduling. In this model, transfer cadence emerges from the interaction between co-designed hardware resources, the host operating system, and local burst buffers, rather than from external queueing policies. As a result, throughput stability is bounded by physical system limits instead of orchestration-induced control latency, enabling predictable performance across heterogeneous deployment environments.
@@
This design avoids the common mismatch between I/O demand and storage capability, which occurs when high-rate network ingress hits latency-dominated production storage services—without requiring application-level awareness or manual tuning.
II-C Integrated Appliances vs Software-Centric Approaches
@@
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving lasting production value after initial deployment. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials [14, 15, 28], and [40]—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations are not equipped to sustain.
The resulting variability makes performance difficult to predict, reproduce, and operationalize across organizations and deployment environments, even when individual components perform well under controlled validation conditions.
@@
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how they can be overcome. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The overall approach is grounded in technology developed over the past decade to meet the ambitious data movement requirements of the premier United States (U.S.) Department of Energy (DOE) Exascale Computing Preparation Project, LCLS-II [53]. The core challenges and performance targets defined by the project directly motivated the architecture. Furthermore, the ongoing dialogue with LCLS-II and other tier-1 scientific entities continues to advance and validate the evolution of this holistic approach to data movement systems. The core software’s performance has been independently validated on long-distance 100 Gbps links, encompassing petabyte-scale transfers with ESnet [14], [15], and [34], SCA19 DMC [40], and transcontinental production trials [28].
@@
The outcomes documented by independent institutions, e.g., [14, 15, 40] are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). This demonstrates that reliable, efficient data movement is most effectively achieved through complete systems engineering [33, 32]. The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing we will note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
@@
Despite the recognized challenges of end-to-end data movement, the prevailing industry and research approach continues to favor complex tuning methodologies and reliance on prohibitively expensive commercial testbeds or bespoke hardware. This complexity is particularly acute at the network edge — the "headwaters" of the “Drainage Basin Pattern”, Fig.1, where resource-constrained environments such as hospitals, clinics, and remote laboratories typically utilize 1–10 Gbps links. Deploying reliable, high-efficiency data transfer in these settings becomes cost-prohibitive when conventional solutions require specialized IT staff and expensive hardware.
@@
The architecture proposed in this paper directly addresses these barriers by validating a different principle: that architectural co-design can inherently reduce both operational complexity and capital expenditure across the entire data spectrum. This principle is practically realized in a mini-appliance designed for resource constrained edge sites. The hardware costs approximately \$2,000 (Fig. 3 Mini Appliance). This demonstrates that high-efficiency 1–10 Gbps data transfer is achievable simply and affordably, without complex system integration or the utilization of inappropriate virtualization options (Section III-F). This cost efficiency and simplicity are a direct consequence of the holistic approach detailed in the following sections.
III Six Common Paradigms
III-A The Latency Killer Paradigm

### v4

# Reexamining Paradigms of End-to-End Data Movement

@@ Abstract.
End-to-End Data Movement, Co-design, Data Movement Appliances, Burst Buffer, System-level Optimization, Cloud, Cross-Domain Data Flow, Research and Education Networks
$c$s: Computer systems organization Distributed architecturesccs: Computer systems organization Heterogeneous hardwareccs: Networks Network performance analysisccs: Information systems Storage management
Preprint
This is a preprint of a research manuscript currently undergoing peer review.
The work introduces a co-design principle for moving data efficiently and the Drainage Basin Pattern (Fig. 1) conceptual model. Successive implementations have undergone over a decade of validation through production-scale deployments from 10 Gbps to 100 Gbps production links, including U.S. Department of Energy (DOE) ESnet technical evaluations (Kissel and Fang, 2020c, b, a) and historical transcontinental production trials (ICM, 2019).
https://creativecommons.org/licenses/by/4.0/
For the latest version, please see:
https://arxiv.org/abs/2512.15028
Authors’ Contact Information:
Chin Fang, Zettar Inc., USA, fangchin@zettar.com
Timothy Stitt, F. Hoffmann-La Roche Ltd., Switzerland
Michael J. McManus, Intel Corporation (Retired), USA
Toshio Moriya, Institute of Materials Structure Science, KEK, Japan
For academic discussion only.
1. Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps (Energy Sciences Network (2025a), ESnet). However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
@@ 1. Introduction
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap manifests as lower than desired data rates, even on 10 Gbps links and modest commodity hardware. Higher-speed networks only amplify its effect. The “Drainage Basin Pattern” conceptual model (Fig. 1) addresses it by shifting the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment). It also enables efficient 1–10 Gbps transfers at the edge. Detailed methodology and measurement results appear in Sections 3 and 4, where we analyze how host configurations influence end-to-end throughput across multiple deployment scenarios. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
@@
Figure 1. The experience of moving data for most practitioners is typically limited to the source of the river. This corresponds to end-user activities, such as transferring photos and videos from mobile phones or moving spreadsheets from one folder to another. At such low data rates, operations appear simple. This may explain why efficient data transfer has historically received limited attention, motivating the introduction of the Drainage Basin Pattern.
A conceptual diagram showing the Drainage Basin Pattern for data movement. Illustrates how data flows from edge sources through networks to core data centers, similar to water flowing from streams into rivers to ocean.
@@
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II (Thayer and Fang, 2019), sustained petabyte transfers (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet), and international deployments (ICM, 2019)—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations (Kissel and Fang, 2020c, b, a), historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN). Together, this body of evidence confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
@@
A key architectural component for achieving predictable performance is the use of high-speed burst buffers (Bhimji et al., 2016) as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) (NSCC, 2019) and official ESnet evaluations (Kissel and Fang, 2020c, b, a), underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
@@
The core contribution of this paper is the identification and validation of six widely held engineering paradigms. They continue to shape expectations, system design, and investment decisions in production end-to-end data movement. Through production-scale deployments and operational trials, we show why they often fail to predict achievable performance across the full edge-to-core spectrum. Reliable and predictable data movement emerges from holistic, co-designed, and tightly integrated hardware-software systems, instead of isolated optimization of individual components. The end-to-end behavior of these systems can be reasoned about using the proposed Drainage Basin Pattern.
@@
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the emphasis placed on standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for premium CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 (Red Hat, 2025b) or free rebuilds such as Rocky Linux 9.6 (Rocky Enterprise Software Foundation, 2025).
2. Background / Related Work
Research programs such as the Global Research Platform (GRP) ((2025a), GRP) have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop ((2025b), GRP) and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem (Fang et al., 2016; Fang, 2019).
@@ 2. Background / Related Work
2.1. Burst Buffers and Data Staging
We now introduce the burst buffer (Bhimji et al., 2016)—a high-speed intermediate storage layer. Originally pioneered in supercomputing to absorb data for high-performance computing (HPC) nodes at extreme velocities, we generalize this concept for sustained, wide-area data movement.
@@ 2.1. Burst Buffers and Data Staging
The burst buffer serves both as a staging area between production storage and the network. A vendor-agnostic data movement appliance design employs a unified software data mover (Zettar zx (Zettar Inc., 2025). See Table 1 for a summary of its features and capabilities) and burst buffers to create a predictable performance envelope that enables sustained near-line-rate data movement between burst buffers (fast ($100~\mathrm{Gbps}$) and high-speed ($\geq 100~\mathrm{Gbps}$) WAN links)
In this architecture, the burst buffer serves both as a fast storage tier and as a deliberate decoupling mechanism. It isolates the data mover from production storage services, which are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors.
@@
2.2. Bulk and Streaming Transfers
@@
  • Streaming Transfer: the movement of a dynamic, actively growing dataset where data is transferred concurrently as it is being generated and written to the source storage. This type is essential for real-time workloads, such as data acquisition from scientific instruments like LCLS-II (Thayer and Fang, 2019) and large-scale bioscience imaging operations (e.g., Cryo-EM, virtual staining) that, due to their inherent data rates, will require such concurrent, line-rate data movement to achieve efficient production workflows.
This distinction is critical because streaming transfers demand both high storage throughput and low storage latency, stressing storage subsystems, memory hierarchies, and scheduling mechanisms in ways that differ fundamentally from bulk transfer. As a result, streaming workloads require architectures designed for sustained concurrency and flow.
@@
The architecture presented in this work eliminates this source of variability by removing centralized orchestration from the data path entirely. Data movement is executed as a decentralized, peer-to-peer service, coordinated implicitly through asynchronous buffer state rather than explicit global scheduling. In this model, transfer cadence emerges from the zx-managed interaction among co-designed hardware resources, the host operating system, and local burst buffers, rather than from external queuing policies. As a result, throughput stability is bounded by physical system limits instead of orchestration-induced control latency, enabling predictable performance across heterogeneous deployment environments.
@@
This design resolves the critical mismatch that occurs when high-rate network ingress meets latency-bound production storage—a problem that lies beyond the expertise of most operational IT staff. It does so without requiring application-level awareness or manual tuning.
2.3. Integrated Appliances vs Software-Centric Approaches
@@
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving target data rates reliably in production. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019)—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations are not equipped to sustain.
The resulting variability makes performance unpredictable and hard to reproduce across different organizations and environments.
@@
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how to do it right. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The overall approach is grounded in technology developed over the past decade to meet the ambitious data movement requirements of the premier United States (U.S.) Department of Energy (DOE) Exascale Computing Preparation Project, LCLS-II (Thayer and Fang, 2019). The core challenges and performance targets defined by the project directly motivated the architecture. Furthermore, the ongoing dialogue with LCLS-II and other tier-1 scientific entities continues to advance and validate the evolution of this holistic approach to data movement systems. The core software’s performance has been independently validated on long-distance 100 Gbps links, encompassing petabyte-scale transfers with ESnet (Energy Sciences Network (2017), ESnet), (Energy Sciences Network (2018), ESnet), and (Kissel and Fang, 2020c), SCA19 DMC (NSCC, 2019), and transcontinental production trials (ICM, 2019).
@@
The outcomes documented by independent institutions, e.g., (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; NSCC, 2019) are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). This demonstrates that reliable, efficient data movement is most effectively achieved through complete systems engineering (Kissel and Fang, 2020b, a). The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing we will note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
@@
Despite the recognized challenges of end-to-end data movement, the prevailing industry and research approach continues to favor complex tuning methodologies and reliance on prohibitively expensive hardware. Such factors are particularly acute at the network edge — the "headwaters" of the “Drainage Basin Pattern”, Fig.1, where resource-constrained environments such as hospitals, clinics, and remote laboratories typically utilize 1–10 Gbps links. Deploying reliable, high-efficiency data transfer in these settings becomes cost-prohibitive when conventional solutions require specialized IT staff and expensive hardware.
@@
The architecture proposed in this paper directly addresses these barriers by validating a different principle: that architectural co-design can inherently reduce both operational complexity and capital expenditure across the entire data spectrum. This principle is practically realized in a mini-appliance designed for resource constrained edge sites. The hardware costs approximately \$2,000 (Fig. 3 Mini Appliance), including high-performance NVMe burst buffers. This demonstrates that high-efficiency 1–10 Gbps data transfer is achievable simply and affordably, without complex system integration or the utilization of inappropriate virtualization options (Section 3.6). This cost efficiency and simplicity are a direct consequence of the holistic approach detailed in the following sections.
3. Six Common Paradigms
3.1. The Latency Killer Paradigm